OpenStudy (anonymous):

if an equation has only an x^2 term how would you solve give an example

OpenStudy (anonymous):

For ax^2+bx+c=0 x = \[(-b \pm \sqrt{b^{2}-4ac}) \over (2a) \]

OpenStudy (anonymous):

So for x^2 =0 x =\[ -0 \pm \sqrt{0^{2}-4*1*0} \over 2*1 \] =0 or 0

OpenStudy (anonymous):

It's useful to get in the habit of using this method however all you really need to do is square root what x-squared equals but the first method works when you have x terms as well

OpenStudy (anonymous):

so what do i write as my answer she wants a complete sentance

OpenStudy (anonymous):

To solve an equation that only has an x^2 term rewrite the equation so you have x^2 on one side of the equation then square root the other side remembering to give the positive and negative root.

OpenStudy (anonymous):

eg \[x^2 -9=0\] \[x^2=9\] \[x= \pm 3 \]